### v1

# Signage-Aware Exploration in Open World using Venue Maps

Abstract
Current exploration methods struggle to search for shops in unknown open-world environments due to a lack of prior knowledge and text recognition capabilities. Venue maps offer valuable information that can aid exploration planning by correlating scene signage with map data. However, the arbitrary shapes and styles of the text on signage, along with multi-view inconsistencies, pose significant challenges for accurate recognition by robots. Additionally, the discrepancies between real-world environments and venue maps hinder the incorporation of text information into planners. This paper introduces a novel signage-aware exploration system to address these challenges, enabling the robot to utilize venue maps effectively. We propose a signage understanding method that accurately detects and recognizes the text on signage using a diffusion-based text instance retrieval method combined with a 2D-to-3D semantic fusion strategy. Furthermore, we design a venue map-guided exploration-exploitation planner that balances exploration in unknown regions using a directional heuristic derived from venue maps with exploitation to get close and adjust orientation for better recognition. Experiments in large-scale shopping malls demonstrate our method’s superior signage recognition accuracy and coverage efficiency, outperforming state-of-the-art scene text spotting methods and traditional exploration methods.
Index Terms:
Scene Understanding, Scene Text Spotting, Exploration-Exploitation, Path Planning.
I Introduction
Searching for shops using their signage information in unknown open-world environments poses a significant challenge for robots due to a lack of prior knowledge. In contrast, humans can effortlessly navigate to the goal landmarks by leveraging the venue maps in daily life. Imagine walking through a shopping mall: we locate ourselves on the venue maps based on actual observations and then plan paths on the maps to reach our destinations. Significant efforts have been made to correlate the visual observations with venue maps for global localization [1, 2, 3, 4, 5], or to discover non-metric heuristic derived from maps for kilometer-scale navigation [6, 7]. However, these approaches primarily focus on using geometric and semantic information from the venue maps, overlooking valuable cues from landmark names on the signage. This oversight is crucial for accurately identifying locations and reducing ambiguity in localization [8]. Moreover, they do not consider online scene understanding and exploration. Current exploration methods generally incorporate geometric information [9, 10] or object-level scene semantics [11, 12, 13, 14, 15, 16], but they often neglect text-level semantics, limiting their effective use of venue maps. While some efforts detect room names as cues in office environments [17, 18, 19], they do not handle signage recognition. We argue that a signage-aware exploration method is essential to bridge the gap between searching for shops using venue maps and exploration methods.
Figure 1: We propose to leverage the venue map as the prior knowledge to facilitate shop searching in unknown open-world environments. In the figure, given a venue map, the robot recognizes the text on the signage of the nearby restaurant and localizes itself on the map. Then the robot knows the direction towards next landmark, which is called ‘Briketenia,’ and searches for it.
However, implementing such a signage-aware exploration method poses two main challenges. First, recognizing shop names on signage in open-world environments is difficult as they are presented in arbitrary shapes and styles, along with multi-view semantics inconsistencies. Existing works in text recognition for visual localization [1, 2] or place recognition [20, 21, 22, 23, 24] often rely on the OCR models trained on closed-set datasets, which struggle to recognize signage effectively in open-world scenarios. To overcome this, an online signage understanding, i.e., text-centric scene understanding method, and an active signage perception policy are essential. Second, using venue maps to guide exploration planning is challenging due to their inconsistent scales and distortions compared to real-world situations. We propose leveraging the directional relationship between landmarks on venue maps as a heuristic to guide exploration without relying on metric information.
This paper presents a signage-aware exploration system that leverages venue maps to facilitate searching for shops in the unknown open world. Our method performs a human-like retrieval process rather than directly employing a conventional closed-set text recognition approach. It identifies the landmark name on the signage by finding the closest match within a pre-established text set extracted from venue maps based on visual similarity. This allows the closed-set detectors to adapt to open-world situations without fine-tuning. We also project multi-view 2D text regions into 3D space and fuse the corresponding features to enhance recognition accuracy. Furthermore, we design a venue map-guided exploration-exploitation planner that enables the robot to search for the signage using a directional heuristic. Once a signage candidate is detected from a distance, the system switches to exploit known spaces to approach the signage and adjust the robot’s view to recognize the signage clearly. By balancing exploration and exploitation, we achieve both high signage recognition accuracy and coverage efficiency. Our main contributions to this work are summarized as follows:
  • A novel signage understanding method comprising a diffusion-based text instance retrieval and a 2D-to-3D semantic fusion strategy that can accurately detect and recognize the signage in the scene.
  • A venue map-guided exploration-exploitation planner for efficiently covering the signage in the open world and meanwhile establishing a signage map.
  • Evaluation experiments in real-world shopping malls have substantiated our method’s efficacy and robustness.
Figure 2: Overall framework of the proposed method. Our method first pre-builds a topological graph based on the spatial relationship between landmark names on a given venue map (Sec. IV-A). Then, given the RGB-D image and point cloud at each timestep, our method first employs a scene text detector to detect the text instances on the observed signage. Next, we project the 2D text regions to 3D space and perform 3D fusion if they can be merged with the previously detected text instances. Then, we introduce a diffusion-based text instance retrieval method to correlate the detected text with the text set on the venue map (Sec. IV-B). Once localized on the venue map, the next landmark goal is inferred by the topological planning and used to guide the selection of frontiers. The detected frontiers and the viewpoints generated at the signage are jointly fed to the planner for a decision on the next best pose. Throughout the process, our approach balances exploration and exploitation to improve signage recognition accuracy and coverage efficiency (Sec. IV-C).
II Related works
II-A Discovering the Potential of Venue Maps
Venue maps, satellite maps or OpenStreetMaps (OSM), are commonly encountered in our daily lives, offering valuable and readily accessible information for human and robot navigation. Existing works primarily study estimating the global pose $(x,y,\theta)$ on the 2D map’s coordinate using visual observation over the latent space [1, 2, 3, 4, 5], such as SNAP [5], which aligns the query image with a fuse of ground-level and overhead imagery over a learned 2D neural space, or inducing the potential heuristics from the map to achieve long-range point-goal navigation in open-world scenarios [6, 7]. ViKiNG [6] learns a heuristic function to predict whether future steps lie in the paths towards the goal, and S2MAT [7] utilizes a user-specified GPS sequence to guide the navigation. As for leveraging the landmark names on the venue maps, [1] introduces the shop names into the global localization and formulate the problem as inference in a Markov random field, and [2] employs a particle filter into text detection for pose estimation. However, correlating shop names with the venue maps to facilitate exploration planning has not been studied.
II-B Semantic-aware Exploration
Semantic-aware exploration methods that focus on the semantics in the scene have arisen in recent years [11, 12, 13, 14]. They integrate scene understanding methods with traditional planning, seeking both mapping efficiency and semantic coverage. SARHP [14] proposes an object-centric planner for optimizing the semantic gain of the exploration paths. 3D scene graph methods that focus on the objects and their topological relations to enable high-level reasoning for robotic navigation and manipulation have also been developed, such as ConceptGraphs [15] and HOV-SG [16]. They segment the object masks via a proposal network such as SAM [25] and extract the language-aligned features of the cropped regions via CLIP [26] to support open-vocabulary query, which assumes a sequence of RGBD trajectory is pre-collected manually or by any exploration method and focus on establishing scene representation on it. Different from the above works that concentrate on object-level semantics, and also distinct from studies targeting simple text cues, e.g., the room names in office environments [17, 18, 19], our work focuses on incorporating the text information on signage into exploration planning.
II-C Text-based Place Recognition and Scene Text Spotting
Text features have garnered significant attention in visual place recognition (VPR) task [20, 21, 22, 23], which employs text-based descriptors to facilitate matching between a query image and a database of scene images, due to their robustness to the appearance changes and distinctive features in different places. Recently, TextSLAM [24] develops a novel SLAM system tightly coupled with text features to build a 3D volumetric map and enhance the pose estimation accuracy. On the other hand, scene text spotting (STS) task that detects and recognizes the text in natural scenes have drawn great attention recently, particularly spotting the text on the signage, which is more difficult in terms of styles and fonts. ESTextSpotter [27] addresses this by explicitly synergying the text detection and recognition modules, achieving state-of-the-art performance.
Despite using the optical character recognition (OCR) models such as PP-OCRv3 [28] and STS models like ESTextSpotter [27], the signage recognition results are often too noisy due to the partial observation and multi-view inconsistency issues. [1] adopts bag-of-K gram to mitigate the former issue, which requires predicting a 10000-dimensional classification vector and is still unsuitable for square characters. To make a general solution, we conduct similarity matching between the observed signage and the text set of venue maps to improve the signage recognition accuracy without fine-tuning. This is similar to the scene text retrieval (STR) task but with different motivations. Prior works in STR convert the queried text to an image via font rendering and perform image-level similarity matching to localize and retrieve all the text instances from a pre-established image gallery [29, 30, 31]. In contrast, we aim to improve the signage recognition accuracy and use a text-diffusion model for text-to-image conversion.
III Problem Formulation
We aim to design an online signage-aware exploration method that leverages the venue map for searching for shops in unknown large-scale human-oriented environments such as shopping malls. We model the environment as a 3D space $V\in\mathbb{R}^{3}$, where the points $\mathbf{v}\in V$ are classified into occupied, free, and unknown with the associated probability $p_{o}(\mathbf{v})$. The environment also contains $N$ static landmarks that can be shops or restaurants identified by their names $\mathcal{T}$. The venue map M (see Fig. 1) portrays all the landmarks and the inaccurate topological relationship $\mathcal{G}$ among them. Note that we do not necessitate any geometric information or semantic layouts since some maps may not contain them or require additional segmentation that limit the generalization of the method. The planner takes as inputs the venue map M, RGB image $o$, depth $d$, and LiDAR points $\mathcal{X}$ at each timestep, outputs next best pose $c\in\mathcal{F}\cup\mathcal{V}$ among the sampled frontiers $f\in\mathcal{F}$ and viewpoints $v\in\mathcal{V}$, while concentrating on the signage in the scene. During the process, a signage map $\mathcal{M}$ is established online, where the recognized signage text $\tau\in\mathcal{T}$ and the corresponding located regions are stored.
IV Methodology
Our method consists of three parts: 1) topological planning on venue maps, 2) signage understanding, and 3) exploration-exploitation planning. The framework is illustrated in Fig. 2.
IV-A Topological Planning on Venue Maps
We first pre-build a topological graph $\mathcal{G}$ on the venue maps whose nodes are the landmark names detected by an OCR model and edges connect two nodes if their Euclidean distance on the map is lower than a threshold $\gamma_{map}$. For the isolated subgraphs, we iteratively merge them at the nearest node pairs. As such, we solve a travel salesman problem (TSP) to obtain a landmark route $g_{1:K}$. During online exploration, our method searches and recognizes the signage in the scene step-by-step to handle long-horizon planning [32].
Figure 3: The pipeline of signage understanding. (a) At the offline stage, we use a text-diffusion model to render the landmark names on the generative images. (b) At the online stage, we project the detected text images to 3D space and fuse the features with previous observations of the same signage. Then we retrieve the most similar offline images compared to the detected images to determine the labels, which are projected onto a signage map. All the text features are extracted by a scene text spotter $\phi$.
IV-B Signage Understanding
We online detect and recognize the signage using the proposed signage understanding method. As such, the robot is globally localized on the venue maps. Meanwhile, we construct a signage map $\mathcal{M}$ for downstream querying and planning, which stores the recognized signage text $\tau$ and the corresponding located regions. In practice, we assign the clustered contours of the online map that is nearest to the signage as its corresponding landmark’s located regions.
Diffusion-based Text Instance Retrieval: Current OCR or STS models are limited by their closed-set training that cannot handle signage recognition with arbitrary shapes, styles, and multi-view inconsistency issues in the open world (see Fig. 6). To address this, we leverage the text set of the venue maps to perform visual similarity matching that enables the closed-set detectors to adapt to the open world without fine-tuning. Specifically, we employ a two-stage process for signage recognition. At the offline stage, we utilize an off-the-shelf multilingual text-diffusion model, AnyText [33], to convert all the landmark names $\mathcal{L}$ to the text-rendered signage images set $\{o^{\prime}\}$. Then, we adopt ESTextSpotter [27], denoted as $\phi$, to detect text regions and only extract the text features before the recognition heads from these offline-generated images without using the recognition results, thereby creating a prior features pool $\mathcal{D}_{f}$. During online exploration, we utilize the same spotter $\phi$ to detect the signage from the current image observation $o_{t}$ and extract its features $f_{t}$. We then calculate the cosine similarity $s_{\phi}(o_{t},o^{\prime}_{j})$ between the detected text features and those of the generative text images to find the most similar pair. In practice, we also enhance the retrieval process by narrowing down the matching scope to a subset of $h$-hop neighbouring to the previous recognized landmark $\tau_{t-1}$ on the venue maps based on the robot’s real-time location, resulting in better accuracy and faster speed:
$s_{\phi}(o_{t},o^{\prime}_{j})=cos(\phi(o_{t}),\phi(o^{\prime}_{j})),\ j\in Neighbour(\mathcal{G},\tau_{t-1},h).$ (1)
As such, the detected scene text $\tau_{t}$ is recognized and matched to the most similar landmark name at the image level, i.e., $\tau_{t}=\mathcal{T}_{arg\max s_{\phi}(o_{t},o^{\prime}_{j})}$, with a recognition score $s_{\phi}(o_{t},o^{\prime}_{j})$.
2D-to-3D Fusion: To mitigate the noisy text recognition from single-view 2D images, we propose adopting a 2D-to-3D instance fusion strategy [34] to enhance recognition robustness using multiple observations. During exploration, the signage map serves as a 3D text bank $\mathcal{B}$ that maintains the merged 3D text instances $R^{3D}_{j}$. At each frame $o_{t}$, we first project the detected 2D text regions $\{r_{i}^{2D},..,r_{k}^{2D}\}$ to 3D regions $\{r_{i}^{3D},..,r_{k}^{3D}\}$ using depth $d_{t}$, pose $P_{t}$ and camera intrinsics $C$. Subsequently, for each new 3D text region $r_{i}^{3D}$ and each existing 3D text instance $R^{3D}_{j}$ in the bank, if their cosine similarity $s(f^{2D}_{i},f^{3D}_{j})$ exceeds a predefined threshold $\gamma_{s}$ and the 3D distance between their centroid points $d(r_{i}^{3D},R_{j}^{3D})$ is lower than a threshold $\gamma_{d}$, then we fuse two instances by aggregating centroid points and averaging over the features:
$R^{3D}_{j}:=R^{3D}_{j}\cup r_{i}^{3D}~{}\text{and}~{}f_{j}^{3D}:=(nf_{j}^{3D}+f_{i}^{2D})/(n+1).$ (2)
We also periodically merge and filter redundant 3D instances in the bank following the above conditions.
IV-C Exploration-Exploitation Planning
Candidate Next Pose: During navigating to the next landmark $g_{k}$, the candidate next poses $c\in V$ are sampled among the set of unvisited frontiers $\mathcal{F}$ and unvisited viewpoints $\mathcal{V}$. The frontiers $f\in\mathcal{F}$ are sampled at the boundaries between known and unknown spaces, which biases the exploration to unknown regions. The viewpoints $v\in\mathcal{V}$ are sampled nearby the signage, allowing the robot to approach the signage and adjust its orientation for better recognition.
Next Best Pose Selection: We tailor the frontier utility $U_{f}$ and viewpoint utility $U_{v}$ for the decision on the candidate selection, respectively. At the initial stage, we induce the robot to explore its surroundings by selecting the frontiers that maximize the information gain $\sum G$ within the camera’s FOV. After at least two landmarks are found, we align the online map with venue map and estimate all the landmark poses $p_{g_{t}}$ at the world coordinate using random sample consensus (RANSAC) [35] and iterative closest points (ICP) [36]. Therefore, we perform the frontier-based informed search using the relative directions between landmarks on the venue maps. To this end, a directional heuristic is calculated to favor selecting the frontiers closest to the direction towards the subgoal $g_{t}$ (see Fig. 4), which is defined as the normalized inner product between the vectors from the current position $p_{t}$ to the frontier $f_{i}$ and the vector from the current position $p_{t}$ to the position of next subgoal $p_{g_{t}}$. Specifically, we first calculate the transformation matrix $T=[R,t]$ and scale $\alpha$, which is updated whenever a new landmark is found. Then, we back-project the landmark positions from the venue map to the world coordinate. As such, the directional heuristic $h(f_{i})$ is calculated as $h(f_{i})=\frac{(p_{t}-f_{i})\cdot(p_{t}-T^{-1}p_{g_{t}})}{\|p_{t}-f_{i}\|*\|p_{t}-T^{-1}p_{g_{t}}\|}$, and thereby the frontier utility $U_{f}(f_{i},p_{t})$ is computed as:
Figure 4: We estimate the position of next landmark by calculating coordinate transformation between online map and venue map for guiding exploration.
$U_{f}(f_{i},p_{t})=\lambda\mu(p_{t},f_{i})h(f_{i})-{\eta}d(p_{t},f_{i}),\ f_{i}\in\mathcal{F},$ (3)
where $d(p_{t},f_{i})$ is the Euclidean distance from the current pose to the candidate pose representing a navigation cost with a weight $\eta$, $\lambda$ is a weight giving more importance to directional heuristic than the navigation cost. The hysteresis gain [9] $\mu(p_{t},f_{i})$, which equals $\mu_{gain}>1$ if frontier points $f_{i}$ lie within a certain radius $\mu_{r}$ of current position $p_{t}$ otherwise equals 1, is adapted to bias selecting frontier points in the surroundings. On the other hand, the viewpoint utility $U_{v}(v_{j},p_{t})$ is defined to favor the exploitation in known spaces whenever a text instance with a certain confidence is detected:
$U_{v}(v_{j},p_{t})=\beta s_{v}(v_{j})-{\eta}d(p_{t},v_{j}),\ v_{j}\in\mathcal{V},$ (4)
where $s_{v}(v_{j})=s_{\phi}(o_{\tau_{j}},o_{\tau_{j}}^{\prime})$ denotes the score that the signage $\tau_{j}$ belonging to $v_{j}$ is matched to the landmarks on venue map, and $d(p_{t},v_{j})$ is the navigation cost with a weight $\eta$. A factor $\beta$ that determines to what extent of scores we should highlight the viewpoint candidates than frontiers is employed to handle the exploration-exploitation dilemma. Intuitively, the priority of exploiting signage originates from the fact that it can help localize the robot on the venue maps and determine the decision on the next best pose. Finally, we obtain the optimal next pose $c_{t+1}$ among unvisited frontiers and viewpoints that maximizes the overall utility $U(c_{j},p_{t})$:
$U(c_{j},p_{t})=I_{f}(c_{j})U_{f}(c_{j},p_{t})+(1-I_{f}(c_{j}))U_{v}(c_{j},p_{t}).$ (5)
where $c_{j}\in\mathcal{F}\cup\mathcal{V}$ and $I_{f}(c_{j})=1\ \text{if}\ c_{j}\in\mathcal{F}\ \text{else}\ 0$. Then we plan a collision-free path to $c_{t+1}$ using the information RRT* planner from Open Motion Planning Library [37] with a fixed replan rate, and follow the route by a local planner from autonomous exploration development environment [^1] https://www.cmu-exploration.com/, which supports dynamic collision avoidance in the scenes. The overall procedure is summarized in algorithm 1.
Algorithm 1 Signage-Aware Exploration using Venue Maps
0:  Venue map $\mathbf{M}$ including landmark names $\mathcal{T}$ and a topological graph $\mathcal{G}$, pseudo text images $\{o^{\prime}\}$, text spotter $\phi$, RGB $o_{t}$, depth $d_{t}$, LiDAR $\mathcal{X}_{t}$
1:  $\mathcal{M}\leftarrow\emptyset$, $\mathcal{F}\leftarrow\emptyset$, $\mathcal{V}\leftarrow\emptyset$
2:  while not done do
3:     $o_{t},d_{t},\mathcal{X}_{t}\leftarrow$ getObservation()
4:     ${\mathcal{M}},f_{t}\leftarrow$ mappingModule($\mathcal{X}_{t}$)
5:     $\mathcal{F}\leftarrow\mathcal{F}\cup f_{t}$
6:     $R^{2D},f^{2D},v_{t}\leftarrow\phi(o_{t})$
7:     if $R^{2D}$ is not None then
8:        $R^{3D},f^{3D}\leftarrow$ Fusion($R^{2D},f^{2D},d_{t},\mathcal{M}$)
9:        $\tau_{t}\leftarrow$ textInstanceRetrieval($f^{3D},\{o^{\prime}\}$)
10:        $g_{k}\leftarrow$ topologicalPlanning($\tau_{t},\mathcal{G}$)
11:        $\mathcal{V}\leftarrow\mathcal{V}\cup v_{t}$
12:        ${\mathcal{M}}\leftarrow$ semanticProjection($\tau_{t},R^{3D}$)
13:     end if
14:     $c_{t+1}\leftarrow$ getNextBestPose($\mathcal{F},\mathcal{V},g_{k}$)
15:     goTo($c_{t+1}$)
16:  end while
17:  return  a signage map $\mathcal{M}$
V Experimental Results
The goals of our experiments are 1) to investigate the signage recognition accuracy of our proposed scene understanding method (Sec. V-C), and 2) to evaluate our exploration-exploitation planner in terms of the average signage coverage rates and exploration time per signage (Sec. V-D).
V-A Experimental Setup
We evaluate the proposed system in two large-scale shopping malls with four and nine shops or restaurants, respectively. We conduct the experiments on the Scout-mini mobile robot platform shown in Fig. 5. The sensor suite comprises a Realsense D455 RGB-D camera and an OS0-128 LiDAR. The computing is performed on an Intel NUC mini PC with Intel Core i9-12900H and an NVIDIA Jetson AGX Orin module with 16GB GPU memory. The WIFI router is used to build a local network for communication among the robot platform, onboard computers and sensors.
V-B Implementation Details
We elaborate the venue maps of the two scenarios via Google My Maps [^2] https://www.google.com/maps/d/ according to the original Google Maps to showcase the landmarks of interest. We adopt FAST-LIO2 [38] for robot pose estimation, a dynamic object removal module [7] for point cloud segmentation, and OctoMap [39] for constructing the 3D volumetric maps simultaneously. We then squeeze the 3D maps into 2D occupancy grid maps. We detect and filter the frontiers by [9], while we generate the viewpoints by heading towards the signage with a certain distance. The hyperparameters used in our method are listed in Table I.
Transformation Estimation: We adopt MAGSAC [35] together with Levenberg-Marquardt refinement to estimate the transformation matrix between the online and venue map by minimizing the sum of the squared error $E(R,t,\alpha)=\frac{1}{N}\sum_{i=0}^{N}||p_{o}-\alpha(Rp_{g}+t)||^{2}$, where $p_{o}$ denotes the centroid of the online detected landmark and $p_{g}$ refers to that on the venue map. We instantiate the above algorithms using the 2D affine transformation method  [^3] https://docs.opencv.org/3.4.1/d9/d0c/group__calib3d.html.
Figure 5: The details of the wheeled vehicle used in real-world experiments.
TABLE I: Hyperparameters
Parameter Value
$h$ num of hop neighbouring to previous landmark 1
FOV horizontal field-of-view of camera 70.5
$\gamma_{map}$ threshold for merging topological graph 150
$\gamma_{s}$ similarity threshold for fusion 0.5
$\gamma_{d}$ distance threshold for fusion 3
$\eta$ weight for navigation cost 0.1
$\lambda$ weight for directional heuristic 5
$\mu_{gain}$ constant in hysteresis gain 5
$\mu_{radius}$ radius in hysteresis gain 10
$\beta$ factor for balancing exploration and exploitation 9
$v_{linear}$ maximum linear speed 0.8
$v_{angular}$ angular speed for turning to active perception 0.1
Dynamic Object Removal: To alleviate the influence of pedestrians during robot exploration, a dynamic object removal approach inherited from [7] is deployed, where the front-end sub-module uses a range-image-based method to coarsely detect the points from dynamic objects, followed by a static mapping framework in the back-end sub-module. Each frame from the LiDAR data stream is effectively separated into dynamic (pedestrians) and static (building structures) scan points, while only the static part is sent into OctoMap [39].
Prompts: We prompt AnyText [33] model to generate the text images by: a logo of a store, with ‘xx’ written on it, which means: {CJK}UTF8gbsn 一个店铺的logo,写着“xx” in Chinese, where xx is the landmark name extracted from venue maps using CnOCR [^4] https://github.com/breezedeus/cnocr.
V-C Signage Recognition Accuracy
To test our signage understanding approach, we evaluate the signage recognition recall rates on a signage images dataset collected in the two scenarios, comparing it with four alternative methods: 1) using CLIP [26] to perform text-to-image similarity matching between landmark names and detected images, 2) using Chinese-CLIP [40] that is fine-tuned on Chinese dataset, 3) using recognition results of an OCR model, ESTextSpotter [27], with Levenshtein distance as the text-level measurement, 4) Retrieval using font: using the latent features extracted from ESTextSpotter and performing image-level retrieval with font-rendered text images [31]. Our method (Retrieval using diffusion) utilizes a text-diffusion model for text-to-image conversion compared to using font. We collect the signage images of 18 shops and restaurants with 5 different views in real-world shopping malls. The examples of signage recognition in open-world scenarios shown in Fig. 6 demonstrate the difficulty. We consider signage to be correctly recognized if the top-$k$ ($k=1,2$) retrieval result equals the ground truth label and the corresponding similarity is higher than a threshold $\gamma_{recall}=0.3$, which is decided to filter the irrelevant text that is neither written on signage nor within the venue maps during exploration.
The results are reported in Table II. It turns out that the signage recognition capabilities of both CLIP and Chinese-CLIP (ViT-B/16) are poor with only 28.9% of recall@1. We attribute it to the observation that CLIPs cannot extract effective text features from signage images that they produce very close similarities when matched with different landmark names and thus cannot filter out the irrelevant texts appearing in the scene. Moreover, the state-of-the-art STS method [27] still struggles with signage recognition, achieving only 52.2% recall rates due to the noisy recognition results. We also show that the most common text measurement, Levenshtein distance, is infeasible for square characters since it does not consider the similarity of strokes inside the characters. By converting texts to the images using font, e.g., Arial, and calculating cosine similarity with signage images [31], the recall rates are improved by 10%. Furthermore, our diffusion-based approach yields an additional 15.6% improvement over font-rending method. We think it is because the generative images are more realistic and have more contextual features, proving that our approach improves the signage recognition capability of a closed-set detector without fine-tuning. Notably, our method’s recall@2 is 10% higher than recall@1, demonstrating its better recognition capability when the search scope is narrowed to neighbouring landmarks. By adopting a 2D-to-3D fusion strategy, our method achieves more robust results with higher similarity scores. Our method can also adapt to recognizing signage in different languages by only fine-tuning the text detectors, since AnyText [33] itself supports rending various languages on the generative images.
Figure 6: Examples of signage recognition. The first column is the ground truth of the texts. In the remaining columns, red boxes highlight the texts of interest, and the words in white boxes are the corresponding recognition results from ESTextSpotter [27]. The noisy results of directly employing a closed-set text spotter along with the interference from numerous text instances that are not presented on the signage make it difficult for accurate signage recognition.
V-D Signage Coverage Efficiency
We compare our signage-aware exploration algorithm with multiple rapidly exploring randomized trees (RRTs)-based exploration [9]. For a fair comparison, we also equip RRTs-based exploration with the proposed signage understanding module. We choose two starting points in two scenarios and conduct three trials at each point for two methods, respectively. We evaluate the signage coverage efficiency by the signage coverage rates and exploration time per signage with the corresponding standard variance, which are reported in Table III. The former refers to the average number of recognized signage of all the signage in the scene, and the latter is calculated as: $\frac{1}{n}\sum T(i)/S(i)$, where $n$ is the number of trials, $T(i)$ and $S(i)$ are the total exploration time until no detected frontiers and signage coverage number of the $i$th trial, respectively. It turns out that the RRTs-based exploration method detects around 1 and 1.67 of 4 signage at two starting points in scenario 1, respectively. For the larger scenario, it can only find 3 and 4 of 9 signage. In contrast, our method strikes a balance between exploration and exploitation, which contributes to improvement in coverage rates that we can cover 3 and 3.67 of 4 signage in scenario 1 and 7 and 6.67 of 9 signage in scenario 2. Furthermore, despite the exploitation costs more time, our method can correlate the recognized signage with venue maps to know the approximate direction towards the next landmark, resulting in reducing more than 1 time the exploration time per signage than the baseline. In consequence, it produces more efficient exploration paths to cover more signage during the progress (see Fig. 7).
TABLE II: Recall rates of signage recognition ($\uparrow$)
Recall@1 Recall@2
CLIP [26] 28.9% 35.6%
Chinese-CLIP [40] 28.9% 45.6%
ESTextSpotter [27] 52.2% 52.2%
Retrieval using font [31] 62.2% 73.3%
Retrieval using diffusion (Ours) 77.8% 87.8%
TABLE III: Signage coverage rates ($\uparrow$) & exploration time per signage ($\downarrow$).
Scenario 1 Starting point 1 Starting point 2
(4 landmarks) Coverage rates Average time ($s$) Coverage rates Average time ($s$)
[9] 1.50 $\pm$ 0.71 / 4 167.55 $\pm$ 25.51 1.67 $\pm$ 0.58 / 4 162.21 $\pm$ 70.67
Ours 3.00 $\pm$ 0.00 / 4 67.89 $\pm$ 10.78 3.67 $\pm$ 0.58 / 4 74.94 $\pm$ 8.70
Scenario 2 Starting point 1 Starting point 2
(9 landmarks) Coverage rates Average time ($s$) Coverage rates Average time ($s$)
[9] 3.33 $\pm$ 0.58 / 9 186.67 $\pm$ 67.79 4.00 $\pm$ 1.00 / 9 171.55 $\pm$ 47.73
Ours 7.00 $\pm$ 1.00 / 9 93.40 $\pm$ 14.23 6.67 $\pm$ 0.58  / 9 120.58 $\pm$ 30.22
Some qualitative examples of the exploration trajectories are illustrated in Fig. 8. We see that RRTs-based exploration in green blindly explores the unknown boundaries, which can only successfully recognize the signage whenever the robot happens to face to the signage with proper orientations. Moreover, it often stops early before fully exploiting the regions and misses some shops at the boundaries of scenarios. In contrast, the trajectories of our method in red perform exploitation behaviour to actively spot the signage in the scene whenever its utility is higher than that of exploration. Once two signage instances are recognized, our method calculates the directional heuristic to induce the robot to reach the next landmarks rapidly and reduce redundant exploration. Besides, our method plans a landmark route on the venue maps by a TSP solver that enables the robot to traverse the signage in the scenarios with less overlapping areas. Videos of real-world experiments are given at https://youtu.be/sFSb8spZ_bo.
Figure 7: The exploration progress of two methods in four scenario settings. The curves stop at the time when the last new signage is covered, after which the exploration may still continue but no more signage is covered.
V-E Ablation study
An important design choice is the weighting parameter for balancing exploration and exploitation. Therefore, we study the influence of the balancing weight by comparing our method with different $\beta=\{3,9,15\}$ in scenario 2 starting from point 1. The higher the weight, the more importance is given to the exploitation. The experimental results are reported in Table IV and the trajectory comparison is illustrated in Fig. 9. We see that when the weight is too low ($\beta=3$), the planner inclines exploring to the frontiers under the guidance of venue maps, but delays or even misses recognizing signage, resulting in lower coverage rates. Although it can finally find out part of missed signage using the directional heuristic, it takes more time and decline the efficiency (see Fig. 9a). In contrast, as $\beta=15$, the planner biases selecting too many viewpoint candidates that it spends more time on perceiving the potential signage nearby the robot and decreases the efficiency to explore further signage (see Fig. 9b). A moderate balance of exploration and exploitation ($\beta=9$) can achieve both higher signage coverage rates and searching efficiency.
TABLE IV: Ablation results of balance weight $\beta$. Higher weight biases more exploitation behaviors than exploration.
Balance weight Coverage rates Average time ($s$)
Scenario 2 starting point 1 $\beta=3$ 6.00 $\pm$ 1.00 / 9 143.17 $\pm$ 20.10
$\beta=9$ 7.00 $\pm$ 1.00 / 9 93.40 $\pm$ 14.23
$\beta=15$ 6.67 $\pm$ 0.58 / 9 123.99 $\pm$ 14.01
Figure 8: Qualitative examples of the exploration trajectories of our method and baselines in four scenario settings, showing that our method produces more efficient and reasonable exploration paths by using the venue maps.
Figure 9: Ablation experiments for evaluating the influence of balancing weight $\beta$. Triangles and circles show the views and poses of robots when exploitation for ablation methods ($\beta=3$ and $15$).
VI Conclusion
We have presented the first signage-aware exploration method that utilizes venue maps to incorporate text-level information for searching shops in open-world environments. By harnessing venue-specific text sets and retrieving the most similar text, our proposed diffusion-based text instance retrieval effectively detects and recognizes signage in the scene. Additionally, a 2D-to-3D semantic fusion strategy is employed to enhance the recognition accuracy. Moreover, we design an exploration-exploitation planner to guide exploration using venue maps while balancing the exploration of new areas with the exploitation of known spaces to recognize signage actively. Experimental results demonstrate that our method surpasses state-of-the-art scene text spotting methods and traditional exploration techniques in terms of both signage recognition accuracy and coverage efficiency. Our method is applicable to navigation tasks by setting goals directly on the venue maps.
A limitation of the proposed method lies in the situations where the text is absent. A promising direction is integrating the structure-aware localization methods [3, 4] to improve exploration efficiency. Another potential work is to enable robots to check if previous signage recognition is wrong when the current signage is not the same as the expected landmark name. Introducing shopfront detection methods [41] can also help find the shops in scenarios with irregular layouts.